# The impact of future D and B meson measurements with the SMOG2 program at LHCb on the determination of nuclear parton distribution functions

### v1

@@ III.3 Pseudo-data generation
where $\sigma_{p{\rm A}}\to c\overline{c}$ is the cross section of the bare charm quark production generated by the MG5aMC code with asymmetric collisions within the SMOG2 experiment rapidity range, and $f(c\to D^{0})=0.542\pm 0.024$ Gladilin:2014tba is the fragmentation factor describing the probability of charm quark to hadronise into $D^{0}$ meson. The factor $2$ is used to include both $D^{0}$ and $\overline{D}^{0}$ production. Note that here and in what follows we take $m_{c}=1.5$ GeV, and we assume that the rapidity of the single charm quark is equal to the rapidity of the meson Cacciari:1996wr. The $D$-meson cross section obtained by MG5aMC describes the SMOG $p$Ne data LHCb:2022cul within the theoretical scale uncertainties, the central values being below the data by a factor of two. The expected yield $N_{D^{0}}$ of observed $D^{0}$ mesons in the experiment is calculated taking into account the overall measurement efficiency, $\varepsilon$, the branching ratio for the decay channel used in the measurement $Br(D^{0}\to K^{-}\pi^{+})=(3.947\pm 0.09\pm 0.12)\cdot 10^{-2}$ ParticleDataGroup:2024cfk, and the sampled luminosity $\mathcal{L}$:
@@
The luminosities for each collision system are provided in Table 1 and are based on Ref. Bursche:2018orf.
@@
where $\sigma_{pA\to b\overline{b}}$ is the cross section of the bare bottom quark production, and the fragmentation factor $f(b\to B^{+})=0.404\pm 0.006$ ParticleDataGroup:2024cfk accounts for the probability of $b$ quark hadronising into a $B^{+}$ meson. Here and below, we use $m_{b}=4.75$ GeV. The expected $N_{B^{+}}$ yield is given by
@@
with $Br(B^{+}\to\overline{D}^{0}\pi^{+})=(4.61\pm 0.10)\cdot 10^{-3}$ ParticleDataGroup:2024cfk and $Br(\overline{D}^{0}\to K^{+}\pi^{-})=(3.947\pm 0.09\pm 0.12)\cdot 10^{-2}$ ParticleDataGroup:2024cfk. The luminosity values are provided in Table 1 (values in parenthesis) and are based on luminosity estimations for a cell target similar to SMOG2 and for a full year of data taking with the proton beam Hadjidakis:2018ifr.
@@
III.4 PDF reweighting
The impact of the SMOG2 pseudo-data on the nPDFs can be studied using the PDF reweighting method Giele:1998gw; Ball:2010gb; Sato:2013ika; Paukkunen:2014zia, which we briefly summarise here. The application of this method was initially used in the context of Monte Carlo PDFs Ball:2010gb which are defined via a set of PDF replicas. In such a case central predictions and the corresponding uncertainties for any PDF-dependent quantity, $\cal{O}$, can be computed as average and variance
@@ III.4 PDF reweighting
Through Bayes theorem bayes1763lii the information encoded in the PDFs can be updated with the additional information provided by new data. On a technical level, in the PDF reweighting, this is done by introducing appropriate weights into Eq. (7) and Eq. (8). Taking into account that we will work with Hessian PDF sets, we can define modified Giele-Keller weights as follows Paukkunen:2014zia; Kusina:2016fxy:
@@
Additionally, to qualitatively assess the efficiency of the reweighting procedure, one can compute the effective number of replicas $N_{\rm eff}$ Ball:2010gb:
@@
In what follows, we will work with Hessian PDFs from nCTEQ Segarra:2020gtj; Duwentaster:2022kpv and EPPS21 Eskola:2021nhw fits. We will follow the procedure of Ref. Kusina:2016fxy: before applying the reweighting procedure, we first convert the Hessian PDF sets (represented by central PDF and a set of error PDFs) into PDF replicas, $f_{k}$, as
@@
where $f_{0}$ is the central PDF set, $f_{i}^{(+)}$ and $f_{i}^{(-)}$ are the plus and minus error PDFs corresponding to the eigenvector direction $i$, $N$ is the number of Hessian eigenvectors and $R_{ki}$ are random numbers from a Gaussian distribution with mean 00 and standard deviation of $1$, which are different for each replica $k$ and each eigen-direction $i$.
For our purposes, we generate $N_{\rm rep}=10^{4}$. Moreover, we note that the nPDFs we adopt in this work are delivered with uncertainties corresponding to a 90% confidence level (c.l.). Since our reweighting procedure aims at obtaining new PDFs at a 68% c.l., the tolerance $T$ in Eq. (9) must be rescaled using the following relation: $T_{68\%}=T_{90\%}/(1.645)^{2}$ Martin:2009iq. In the case of nCTEQ, $T_{90\%}=35$, while in the case of EPPS21, $T_{90\%}=33$. For our purposes, as the two tolerances are very close to each other, we adopt for simplicity $T_{90\%}=35$, resulting in a rescaled tolerance $T_{68\%}\approx 13$.
@@ IV Results
Figure 2: $R_{pA}$ predictions for $D^{0}$ production in proton-nucleus collisions for different nuclei (He, Ne, Ar, Kr, Xe) as a function of the atomic mass number for each considered $y_{\rm cms}^{D^{0}}$ interval within the LHCb detector acceptance. The nPDF uncertainties from nCTEQ15HQ Duwentaster:2022kpv are shown with blue boxes.
@@
We begin by demonstrating the current nPDF uncertainties for $D^{0}$-meson production process. In the kinematic region of interest, the scale uncertainty for $R_{pA}$ nearly cancels between numerator and denominator, and is negligible compared to the PDF errors. Hence, in what follows, we will only concentrate on the nPDF uncertainty and we will not vary the scale. Figure 2 shows the nuclear modification factors $R_{pA}$ for He, Ne, Ar, Kr and Xe nuclei for various $D^{0}$-meson rapidity intervals. The boxes represent the current nPDF uncertainties (we adopt here the nCTEQ15HQ nPDFs Duwentaster:2022kpv as an example). They are the largest for the lowest values of rapidity, which corresponds to relatively large $x$, which corresponds to the end of the EMC region and the beginning of the antishadowing region. Clearly, the uncertainties due to nPDFs are far from being satisfactory. It can be also seen from Fig. 2 that the uncertainties are larger for heavier nuclei: this is partially driven by the nPDF parameterisations which typically underestimate the uncertainties for light nuclei.
@@
We will use the pseudo-data generated according to the procedure presented in Section III.3 and the PDF reweighting method described in Section III.4 to see the potential constraints that can be provided by the future LHCb SMOG2 measurement of $D^{0}$ and $B$ meson on the current nPDFs. For this purpose, we use a selection of recent nPDFs: EPPS21 Eskola:2021nhw, nCTEQ15HQ Duwentaster:2022kpv, and nCTEQ15HIX Segarra:2020gtj. However, before presenting the numerical results, we would like to point out some caveats of such an analysis, which prevents us from showing the full potential of the fixed-target LHCb data. First of all, as already said earlier the uncertainties of gluon nPDFs at large-$x$ are underestimated. This is caused by the combination of two issues: (i) a small amount of available data that allows to constrain this region, and (ii) the limits of the Hessian method, used to obtain nPDF uncertainties, to provide realistic error estimates in the absence of data constraints. Furthermore, in most cases the uncertainties for light-nuclei are systematically underestimated due to the way the nPDF parameterisations are constructed. Finally, what is probably the most important limitation is that in order to produce pseudo-data for this study we need to base them on theoretical predictions using nPDFs. Typically, this is not a problem, as the crucial part is to have realistic estimates of the uncertainties of pseudo-data. However, in the current study this also means that the pseudo-data we will use encodes the nuclear $A$-dependence of the corresponding nPDFs. Hence, our analysis has limited capability to study the $A$-dependence of nPDFs. Nevertheless, we would like to highlight that the knowledge of the $A$-dependence is very limited, therefore the future SMOG2 measurements will give a unique opportunity to determine it and to test the assumptions that are currently used in nPDF fits to model it.
@@
Figure 3 shows $R_{pA}$ as a function of the rapidity of the $D^{0}$ meson for the corresponding nuclei, both before and after the reweighting using the two scenarios. We use here the EPPS21 nPDFs as a representative example, while later we provide additional comparisons with results obtained with different nPDFs. The SMOG2 pseudo-data cover midrapidity to moderate backward rapidity region in the cms frame, which in the fixed-target mode is the region where the antishadowing effect is expected (the corresponding $x$ values can be read from Fig. 1c). This fact is reflected in Fig. 2 where the values of $R_{pA}$ for all rapidity intervals are higher than unity.
Figure 3: $R_{pA}$ predictions for the $D^{0}$-meson production in $pA$ collisions for different nuclei (He, Ne, Ar, Kr, Xe) for the EPPS21 nPDFs Eskola:2021nhw. Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting for scenarios 1 and 2 are shown with grey, blue and yellow bands respectively.
Figure 4: $R_{p{\rm Xe}}$ predictions for the $D^{0}$-meson production in $p$Xe collisions for different sets of nPDFs (nCTEQ15HQ Duwentaster:2022kpv, nCTEQ15HIX Segarra:2020gtj, EPPS21 Eskola:2021nhw). Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting are shown with grey, blue and yellow bands respectively.
@@
In Fig. 5 we present the results of the reweighting using $R_{pA}$ as a function of the rapidity of the $B^{+}$ meson in $p$Xe collisions for the same set of nPDFs. Pseudo-data are shown in black, and nPDF uncertainties before and after the reweighting are shown as grey and blue bands, respectively. In this case, we restrict the reweighting to the pseudo-data for Xe only (corresponding to scenario 1 for the $D$ meson). Notice that the assumed luminosities for $B$ mesons are larger than the ones for $D$ mesons (see Tab. 1 and Section II.2). [^3] For the $B$-meson measurements, uncertainties will be limited by the statistics, whereas for the $D$ meson they are expected to be limited by the systematic uncertainties. The results exhibit similar features as for the $D$ mesons, as we observe a reduction of uncertainties in the large negative rapidity region. However, in the case of $B^{+}$ mesons, the probed $x$ values are larger, getting closer to the EMC region (we indeed start to see suppression in $R_{p\mathrm{Xe}}$). This is something that is reflected in the results for the nCTEQ15HIX nPDFs. Here the uncertainties at larger negative rapidities are reduced more than in the case of $D$ meson (see Fig. 4), and the reduction is more prominent when the rapidity is more negative. Even though the probed rapidity range and the expected measurement accuracy are limited, the $B^{+}$ data will allow a significant improvement in our description of the EMC effect for gluons.
Figure 5: $R_{p{\rm Xe}}$ predictions for $B^{+}$-meson production in $p$Xe collisions for different sets of nPDFs (nCTEQ15HQ Duwentaster:2022kpv, nCTEQ15HIX Segarra:2020gtj, EPPS21 Eskola:2021nhw). Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting are shown with grey and blue bands respectively.
Figure 6: Gluon nPDF for Xe nucleus for nCTEQ15HQ Duwentaster:2022kpv and EPPS21 Eskola:2021nhw at $Q=3$ GeV and $Q=9.5$ GeV.
Figure 7: Gluon nPDF for Pb nucleus for nCTEQ15HQ Duwentaster:2022kpv and EPPS21 Eskola:2021nhw at $Q=3$ GeV and $Q=9.5$ GeV.
Figure 8: Strange-, anti-up-, and anti-down-quark distributions in Pb nucleus from the reweighting with nCTEQ15HIX nPDFs Segarra:2020gtj at $Q=3$ GeV.
@@
V Conclusions
In this work we adopted the PDF reweighting method to estimate the potential impact of $D$- and $B$-meson production measurements at the fixed-target program, SMOG2, at LHCb experiment on nuclear PDFs (nPDFs). We have shown that such a program provides an excellent opportunity to improve our understanding of gluon nPDF and that it can even provide some constraints for light-quarks nPDFs. The data expected to be collected at SMOG2 have the potential to significantly enhance the precision of gluon nPDF determination, especially at relatively large longitudinal momentum fraction $x$, where the antishadowing and EMC effects are expected ($0.02\lesssim x\lesssim 0.8$). This region is poorly known, which is reflected in significant uncertainties of nPDFs. At the same time, it is crucial for constraining hadron structure models, learning the properties of QCD confinement, or explaining the EMC effect Higinbotham:2013hta; Malace:2014uea.

### v2

@@ 3.3 Pseudo-data generation
where $\sigma_{pA}\to c\overline{c}$ is the cross section of the bare charm quark production generated by the MG5aMC code with asymmetric collisions within the SMOG2 experiment rapidity range, and $f(c\to D^{0})=0.542\pm 0.024$ [48] is the fragmentation factor describing the probability of charm quark to hadronise into $D^{0}$ meson. The factor $2$ is used to include both $D^{0}$ and $\overline{D}^{0}$ production. Note that here and in what follows we take $m_{c}=1.5$ GeV, and we assume that the rapidity of the single charm quark is equal to the rapidity of the meson [49]. The $D$-meson cross section obtained by MG5aMC describes the SMOG $p$Ne data [31] within the theoretical scale uncertainties, the central values being below the data by a factor of two. The expected yield $N_{D^{0}}$ of observed $D^{0}$ mesons in the experiment is calculated taking into account the overall measurement efficiency, $\varepsilon$, the branching ratio for the decay channel used in the measurement $Br(D^{0}\to K^{-}\pi^{+})=(3.947\pm 0.09\pm 0.12)\cdot 10^{-2}$ [50], and the sampled luminosity $\mathcal{L}$:
@@
The luminosities for each collision system are provided in Table 1 and are based on Ref. [1].
@@
where $\sigma_{pA\to b\overline{b}}$ is the cross section of the bare bottom quark production, and the fragmentation factor $f(b\to B^{+})=0.404\pm 0.006$ [50] accounts for the probability of $b$ quark hadronising into a $B^{+}$ meson. Here and below, we use $m_{b}=4.75$ GeV. The expected $N_{B^{+}}$ yield is given by
@@
with $Br(B^{+}\to\overline{D}^{0}\pi^{+})=(4.61\pm 0.10)\cdot 10^{-3}$ [50] and $Br(\overline{D}^{0}\to K^{+}\pi^{-})=(3.947\pm 0.09\pm 0.12)\cdot 10^{-2}$ [50]. The luminosity values are provided in Table 1 (values in parenthesis) and are based on luminosity estimations for a cell target similar to SMOG2 and for a full year of data taking with the proton beam [23].
@@
3.4 PDF reweighting
The impact of the SMOG2 pseudo-data on the nPDFs can be studied using the PDF reweighting method [51, 52, 53, 54], which we briefly summarise here. The application of this method was initially used in the context of Monte Carlo PDFs [52] which are defined via a set of PDF replicas. In such a case central predictions and the corresponding uncertainties for any PDF-dependent quantity, $\cal{O}$, can be computed as average and variance
@@ 3.4 PDF reweighting
Through Bayes theorem [55] the information encoded in the PDFs can be updated with the additional information provided by new data. On a technical level, in the PDF reweighting, this is done by introducing appropriate weights into Eq. (7) and Eq. (8). Taking into account that we will work with Hessian PDF sets, we can define modified Giele-Keller weights as follows [54, 56]:
@@
Additionally, to qualitatively assess the efficiency of the reweighting procedure, one can compute the effective number of replicas $N_{\rm eff}$ [52]:
@@
In what follows, we will work with Hessian PDFs from nCTEQ [43, 44] and EPPS21 [45] fits. We will follow the procedure of Ref. [56]: before applying the reweighting procedure, we first convert the Hessian PDF sets (represented by central PDF and a set of error PDFs) into PDF replicas, $f_{k}$, as
@@
where $f_{0}$ is the central PDF set, $f_{i}^{(+)}$ and $f_{i}^{(-)}$ are the plus and minus error PDFs corresponding to the eigenvector direction $i$, $N$ is the number of Hessian eigenvectors and $R_{ki}$ are random numbers from a Gaussian distribution with mean 00 and standard deviation of $1$, which are different for each replica $k$ and each eigen-direction $i$ [^4] Such an approach provides a perfect reproduction of PDF errors for Hessian PDF sets using the single tolerance criterion and a symmetric definition of uncertainties [56]. This is the case for the PDFs from the nCTEQ group. In the case of EPPS21 PDFs, one could use a slightly modified prescription provided in [57]. However, since the difference with the asymmetric uncertainty definition is small, we restrict to using the same procedure as in Eq. (14) (which is equivalent to the Hessian reweighting method employed by the EPPS group [54]).
For our purposes, we generate $N_{\rm rep}=10^{4}$. Moreover, we note that the nPDFs we adopt in this work are delivered with uncertainties corresponding to a 90% confidence level (c.l.). Since our reweighting procedure aims at obtaining new PDFs at a 68% c.l., the tolerance $T$ in Eq. (9) must be rescaled using the following relation: $T_{68\%}=T_{90\%}/(1.645)^{2}$ [58]. In the case of nCTEQ, $T_{90\%}=35$, while in the case of EPPS21, $T_{90\%}=33$. For our purposes, as the two tolerances are very close to each other, we adopt for simplicity $T_{90\%}=35$, resulting in a rescaled tolerance $T_{68\%}\approx 13$.
@@ 4 Results
Figure 2: $R_{pA}$ predictions for $D^{0}$ production in proton-nucleus collisions for different nuclei (He, Ne, Ar, Kr, Xe) as a function of the atomic mass number for each considered $y_{\rm cms}^{D^{0}}$ interval within the LHCb detector acceptance. The nPDF uncertainties from nCTEQ15HQ [44] are shown with blue boxes.
@@
We begin by demonstrating the current nPDF uncertainties for $D^{0}$-meson production process. In the kinematic region of interest, the scale uncertainty for $R_{pA}$ nearly cancels between numerator and denominator, and is negligible compared to the PDF errors. Hence, in what follows, we will only concentrate on the nPDF uncertainty and we will not vary the scale. Figure 2 shows the nuclear modification factors $R_{pA}$ for He, Ne, Ar, Kr and Xe nuclei for various $D^{0}$-meson rapidity intervals. The boxes represent the current nPDF uncertainties (we adopt here the nCTEQ15HQ nPDFs [44] as an example). They are the largest for the lowest values of rapidity, which corresponds to relatively large $x$, which corresponds to the end of the EMC region and the beginning of the antishadowing region. Clearly, the uncertainties due to nPDFs are far from being satisfactory. It can be also seen from Fig. 2 that the uncertainties are larger for heavier nuclei: this is partially driven by the nPDF parameterisations which typically underestimate the uncertainties for light nuclei.
@@
We will use the pseudo-data generated according to the procedure presented in Section 3.3 and the PDF reweighting method described in Section 3.4 to see the potential constraints that can be provided by the future LHCb SMOG2 measurement of $D^{0}$ and $B$ meson on the current nPDFs. For this purpose, we use a selection of recent nPDFs: EPPS21 [45], nCTEQ15HQ [44], and nCTEQ15HIX [43]. However, before presenting the numerical results, we would like to point out some caveats of such an analysis, which prevents us from showing the full potential of the fixed-target LHCb data. First of all, as already said earlier the uncertainties of gluon nPDFs at large-$x$ are underestimated. This is caused by the combination of two issues: (i) a small amount of available data that allows to constrain this region, and (ii) the limits of the Hessian method, used to obtain nPDF uncertainties, to provide realistic error estimates in the absence of data constraints. Furthermore, in most cases the uncertainties for light-nuclei are systematically underestimated due to the way the nPDF parameterisations are constructed. Finally, what is probably the most important limitation is that in order to produce pseudo-data for this study we need to base them on theoretical predictions using nPDFs. Typically, this is not a problem, as the crucial part is to have realistic estimates of the uncertainties of pseudo-data. However, in the current study this also means that the pseudo-data we will use encodes the nuclear $A$-dependence of the corresponding nPDFs. Hence, our analysis has limited capability to study the $A$-dependence of nPDFs. Nevertheless, we would like to highlight that the knowledge of the $A$-dependence is very limited, therefore the future SMOG2 measurements will give a unique opportunity to determine it and to test the assumptions that are currently used in nPDF fits to model it.
@@
Figure 3: $R_{pA}$ predictions for the $D^{0}$-meson production in $p$A collisions for different nuclei (He, Ne, Ar, Kr, Xe) for the EPPS21 nPDFs [45]. Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting for scenarios 1 and 2 are shown with grey, blue and yellow bands respectively.
Figure 4: $R_{p{\rm Xe}}$ predictions for the $D^{0}$-meson production in $p$Xe collisions for different sets of nPDFs (nCTEQ15HQ [44], nCTEQ15HIX [43], EPPS21 [45]). Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting are shown with grey, blue and yellow bands respectively.
Figure 3 shows $R_{pA}$ as a function of the rapidity of the $D^{0}$ meson for the corresponding nuclei, both before and after the reweighting using the two scenarios. We adopt a conservative approach and assume uncorrelated systematic uncertainties between the $p$H and $p$A data samples to compute the $R_{pA}$ pseudo-data uncertainties. We use here the EPPS21 nPDFs as a representative example, while later we provide additional comparisons with results obtained with different nPDFs. The SMOG2 pseudo-data cover midrapidity to moderate backward rapidity region in the cms frame, which in the fixed-target mode is the region where the antishadowing effect is expected (the corresponding $x$ values can be read from Fig. 1c). This fact is reflected in Fig. 2 where the values of $R_{pA}$ for all rapidity intervals are higher than unity.
@@
In Fig. 5 we present the results of the reweighting using $R_{pA}$ as a function of the rapidity of the $B^{+}$ meson in $p$Xe collisions for the same set of nPDFs. Pseudo-data are shown in black, and nPDF uncertainties before and after the reweighting are shown as grey and blue bands, respectively. In this case, we restrict the reweighting to the pseudo-data for Xe only (corresponding to scenario 1 for the $D$ meson). Notice that the assumed luminosities for $B$ mesons are larger than the ones for $D$ mesons (see Tab. 1 and Section 2.2). [^5] For the $B$-meson measurements, uncertainties will be limited by the statistics, whereas for the $D$ meson they are expected to be limited by the systematic uncertainties. The results exhibit similar features as for the $D$ mesons, as we observe a reduction of uncertainties in the large negative rapidity region. However, in the case of $B^{+}$ mesons, the probed $x$ values are larger, getting closer to the EMC region (we indeed start to see suppression in $R_{p\mathrm{Xe}}$). This is something that is reflected in the results for the nCTEQ15HIX nPDFs. Here the uncertainties at larger negative rapidities are reduced more than in the case of $D$ meson (see Fig. 4), and the reduction is more prominent when the rapidity is more negative. Even though the probed rapidity range and the expected measurement accuracy are limited, the $B^{+}$ data will allow a significant improvement in our description of the EMC effect for gluons.
Figure 5: $R_{p{\rm Xe}}$ predictions for $B^{+}$-meson production in $p$Xe collisions for different sets of nPDFs (nCTEQ15HQ [44], nCTEQ15HIX [43], EPPS21 [45]). Pseudo-data are shown in black, while the nPDF uncertainties before and after the reweighting are shown with grey and blue bands respectively.
Figure 6: Gluon nPDF for Xe nucleus for nCTEQ15HQ [44] and EPPS21 [45] at $Q=3$ GeV and $Q=9.5$ GeV.
Figure 7: Gluon nPDF for Pb nucleus for nCTEQ15HQ [44] and EPPS21 [45] at $Q=3$ GeV and $Q=9.5$ GeV.
Figure 8: Strange-, anti-up-, and anti-down-quark distributions in Pb nucleus from the reweighting with nCTEQ15HIX nPDFs [43] at $Q=3$ GeV.
@@
5 Conclusions
In this work we adopted the PDF reweighting method to estimate the potential impact of $D$- and $B$-meson production measurements at the fixed-target program, SMOG2, at LHCb experiment on nuclear PDFs (nPDFs). We have shown that such a program provides an excellent opportunity to improve our understanding of gluon nPDF and that it can even provide some constraints for light-quarks nPDFs. The data expected to be collected at SMOG2 have the potential to significantly enhance the precision of gluon nPDF determination, especially at relatively large longitudinal momentum fraction $x$, where the antishadowing and EMC effects are expected ($0.02\lesssim x\lesssim 0.8$). This region is poorly known, which is reflected in significant uncertainties of nPDFs. At the same time, it is crucial for constraining hadron structure models, learning the properties of QCD confinement, or explaining the EMC effect [59, 60].